alones (alones):

Wait..how do I draw equation in slope-intercept 3x - 2y = 5 ?? I actually need to see if I'm right

OpenStudy (anonymous):

What exactly are you asking?

alones (alones):

I'm actually asking..wait so i solved it and i got this solution 

OpenStudy (dariusx):

if x = 0, then -2y = 5 y = -5/2 = -2.5 place one dot on the graph (0;-2.5) then take y = 0 and you will get 3x = 5 x = 5/3 = 1.66 place second dot on the graph (1.66;0) then connect the two dots

OpenStudy (anonymous):

Do you mean write out the equation? Or draw the graph?

alones (alones):

I need just to dri it in a slope intercept form :/

OpenStudy (anonymous):

To put it into slope-intercept form is quite simple. Just isolate and solve for y

alones (alones):

yeah i did so am i right??

OpenStudy (dariusx):

you can choose any number you want for both x and y, but only one at a time take x for example x = 1, then you will get -2y = 5-3 -2y=2 y=-1 and there you have a dot (1;-1)

alones (alones):

xD thx but how would i write it out??

OpenStudy (anonymous):

\[3x-2y=5\] \[-2y=-3x+5\] \[y=\frac{ 3 }{ 2 }x - \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

You are right, but you forgot the x in the one you wrote out above.

alones (alones):

xD o thank you very mucj :)

OpenStudy (anonymous):

@DariusX The question isn't on how to graph but how to put into slope-inercept form. @AloneS No problem.

OpenStudy (dariusx):

oh, sorry slope-intercept form looks like y = ax + b then 3x - 2y = 5 -2y = 5 - 3x y = -5/2 + 3/2x y = 1.5x - 2.5

OpenStudy (dariusx):

nvm i see you already got the answer

alones (alones):

thx guys :)

OpenStudy (anonymous):

@AloneS I can only reply to your messages if you fan me

alones (alones):

O_o shoot i forgot about that :/ sorry

OpenStudy (anonymous):

It's cool. Either fan me or just tell me what you need here :)